# Physics-Informed Transformer operator for the prediction of three-dimensional turbulence

### v3

@@ Abstract
preprint: AIP/123-QED
I Introduction
Turbulence is ubiquitous in both natural phenomena and engineering applications[60]. Due to the complex multiscale nature of turbulence, traditional computational fluid dynamics (CFD) methods face trade-offs between computational cost and accuracy.[51, 64, 65] Various machine learning (ML) methods have been developed to improve or replace CFD simulations of turbulence.[3, 29] Most early machine learning methods were primarily used to enhance the accuracy of turbulence closure models for Reynolds-averaged Navier-Stokes simulation (RANS) and large-eddy simulation (LES).[2, 84, 32, 49, 100, 43, 81, 99, 54, 89].
@@ I Introduction
In recent years, deep neural network methods have been developed for directly simulating flow fields.[66, 63, 52, 4, 62] Raissi et al. proposed physics-informed neural networks (PINNs), a method that embeds physical laws directly into the loss function to address forward and inverse partial differential equations (PDE) problems.[61] Wang et al. proposed the turbulent flow net (TF-Net), a physics-informed architecture designed to capture the dynamics of two-dimensional (2D) Rayleigh-Bénard convection.[72] Jin et al. developed NSFnets, a PINN-based framework that predicts flow fields given their initial and boundary conditions.[27] Although PINNs have demonstrated the capability to achieve satisfactory results even with sparse or no labeled data by embedding physical laws into the loss function, their application in complex scenarios is a great challenge.[94, 78] Wang et al. proposed an adaptive weighting algorithm to mitigate gradient pathologies in PINNs by dynamically balancing the gradient magnitudes of different loss terms.[76] Jagtap et al. proposed adaptive activation functions to accelerate the convergence and improve the accuracy of PINNs by dynamically adjusting the slopes of neurons during training.[25] Yang et al. introduced Bayesian physics-informed neural networks (B-PINNs), which combine Bayesian inference with PINNs to enable stable predictions and uncertainty quantification of physical fields using noisy experimental data.[88] Wang et al. achieved the simulations of 3D turbulent flows by improving PINN training strategies.[74] Wang et al. introduced ShampoO with Adam in the Preconditioner’s eigenbasis (SOAP), a quasi-Newton method that resolves gradient conflicts through second-order optimization, achieving breakthrough accuracy in turbulence simulations at high Reynolds numbers.[73] Zhang et al. used coordinate transformation to mitigate large gradients, enabling accurate PINN simulations of high-Reynolds-number flat plate and NACA0012 airfoil flows[95], while Cao et al. utilized mesh transformation to resolve the sharp leading-edge transitions in subsonic airfoil flows.[7] Moreover, Cao et al. stabilized PINN training by introducing a Jacobian-based gain-tuning strategy, enabling the robust simulation of complex 3D flows.[8] To improve the temporal modeling of PINNs, Zhao et al. proposed Pinnsformer, which leverages Transformer blocks to better capture temporal dependencies than conventional MLPs.[97].
@@
Previous studies primarily focused on neural networks for learning finite-dimensional mappings, often struggling to generalize to other flow types or boundary conditions. To address this limitation, the framework of neural operators (NO) was proposed to learn operators between infinite-dimensional function spaces.[47, 40, 34, 48, 57, 22] Among these, Li et al. proposed the Fourier neural operator (FNO), which approximates the solution operator by parameterizing the integral kernel in the Fourier domain. In 2D Navier-Stokes turbulence, FNO outperformed convolutional networks in terms of accuracy and efficiency.[39] Consequently, numerous extensions of FNO were proposed. You et al. proposed the implicit Fourier neural operator (IFNO), which significantly reduces memory requirements and the number of trainable parameters.[90] Tran et al. proposed the factorized Fourier neural operator (F-FNO), which achieves more efficient operator learning by performing low-rank factorization on the frequency-domain weight tensor.[68] Additionally, Peng et al. embedded an attention mechanism into the FNO model, enhancing its predictive accuracy for 2D turbulence.[58] While early FNO-based studies primarily focused on one-dimensional (1D) and 2D problems, Li et al. extended the framework to 3D space and applied it to the predictions of 3D homogeneous isotropic turbulence.[35] To further address the long-term instability issue in FNO, Li et al. subsequently developed an implicit U-Net enhanced FNO (IUFNO), achieving stable and accurate long-term predictions of 3D turbulent flows.[36] Wang et al. extended IUFNO to 3D turbulent channel flows, achieving fast, accurate and stable long-term predictions.[80]
@@
Although learning parameters in the frequency domain enables FNO to predict turbulence effectively and accurately, it should be noted that FNO relies on the Fourier transform and is therefore primarily applicable to periodic boundaries.[38] Recently, Transformer models have demonstrated outstanding performance in computer vision and natural language processing.[69, 30, 45, 13] Similar to FNO, Transformer also has the ability to capture global information and is increasingly being employed to solve partial differential equations[82, 21, 10] and perform super-resolution flow reconstruction.[71, 79, 98, 18, 33] Cao et al. established a theoretical connection between self-attention and Petrov-Galerkin projection, and proposed Galerkin Transformer to enhance approximation efficiency in PDE problems.[6] Patil et al. proposed a convolutional Transformer model, achieving accurate long-term predictions for 2D wake flows.[56] Xu et al. proposed a super-resolution Transformer (SRTT) to recover fine-scale structures of turbulent flows from low-resolution data.[85]
@@
Despite these successes, the high computational cost of standard attention remains a challenge for 3D simulations. To address this issue, Niu et al. proposed a 3D spatiotemporal model incorporating a channel attention mechanism to predict unsteady airflow over complex three-dimensional aircraft structures.[53] Li et al. proposed the factorized Transformer (FactFormer) for 3D PDE simulations via axial factorization.[37] Yang et al. achieved stable long-term predictions of 3D turbulence by proposing the implicit factorized Transformer (IFactFormer).[86] Yang et al. subsequently improved the model by replacing serial attention with a parallel mechanism, which achieves more accurate long-term predictions of 3D turbulent channel flows.[87]
@@
Another effective approach is the vision Transformer (ViT), proposed by Dosovitskiy et al., which reduces computational costs while achieving excellent results in computer vision by partitioning images into patches.[13] Recently, the ViT has been widely applied to fluid mechanics and aerodynamics problems.[50, 28, 67, 11, 44, 92, 24, 83] Deng et al. proposed a modified ViT-based network for the rapid and accurate predictions of transonic flow fields over supercritical airfoils.[12] Ovadia et al. proposed the vision Transformer neural operator (ViTO), which exhibits an excellent predictive accuracy in 2D partial differential equations.[55] Wang et al. proposed the continuous vision Transformer (CViT) for 2D PDEs.[75]
@@
However, purely data-driven neural operators often suffer from reliance on high-fidelity training data and a lack of physical interpretability. Recently, the physics-informed method has been introduced into neural operators to solve PDEs, including physics-informed DeepONets[77] and physics-informed Fourier neural operator.[41] It has been shown that physics-informed neural operator (PINO) models perform with high accuracy across various differential equations.[20, 42, 26, 9, 70, 14, 16, 93] By leveraging equation constraints, PINO can learn partial differential equations with minimal or even no data. Zhao et al. proposed the large-eddy simulation nets model (LESnets), which performs excellently in predictions of 3D free shear turbulence and isotropic decaying turbulence without using any labeled data.[96] Recently, Lorsung et al. proposed the physics informed token Transformer (PITT) to solve 1D and 2D PDEs.[46] Due to the increase in sequence length, applying Transformer methods to 3D problems requires significant computational costs. Consequently, we consider the ViT method as an effective solution to handle the computational complexity of 3D problems.
@@
II Methodology
II.1 Governing equations of large eddy simulation
The Navier-Stokes equations governing incompressible fluid motion are given by:[60]
@@ II.1 Governing equations of large eddy simulation
Due to the multi-scale nature of turbulence at high Reynolds number, directly resolving all spatial scales requires extremely fine grids, resulting in computational costs that are expensive for practical engineering applications. The LES method applies a spatial filter to the Navier-Stokes equations, enabling the use of coarser grids to resolve large-scale turbulent structures while modeling the unresolved small-scale motions. The filtered governing equations can be written as:[60]
@@
A widely used approach to close the SGS stress is the Smagorinsky (SM) model, which is a representative functional model based on the eddy-viscosity hypothesis.[65, 19] The SM model assumes that the SGS stress is proportional to the resolved strain rate tensor $\bar{S}_{ij}$, formulated as:
@@
where $C_{\text{smag}}$ is the Smagorinsky coefficient and $\Delta$ is the filter width. $\bar{S}_{ij}=\frac{1}{2}\left(\frac{\partial\bar{u}i}{\partial x_{j}}+\frac{\partial\bar{u}_{j}}{\partial x_{i}}\right)$ denotes the filtered rate-of-strain tensor. $|\bar{S}|=\sqrt{2\bar{S}_{ij}\bar{S}_{ij}}$ represents the characteristic strain rate.
The integral length scale and the large-eddy turnover time are defined as:[60]
@@
II.2 Vision Transformer operator
In recent years, Transformer[69] based on self-attention has become the preferred models in natural language processing (NLP). However, it is computationally prohibitive to apply self-attention directly to images due to the quadratic cost associated with the number of pixels. To address this issue, Dosovitskiy et al. proposed the vision Transformer (ViT).[13] ViT reduces the sequence length by partitioning images into patches, thereby reducing the computational overhead of self-attention. After extensive data pre-training, it achieved superior results compared to convolutional neural networks (CNNs).
@@ II.2 Vision Transformer operator
To approximate the nonlinear solutions operator $\mathcal{G}$, the network is constructed as an iterative architecture. Each layer updates the feature field $v_{l}$ via a kernel integral operator $\mathcal{K}$, defined generally as:[96, 36, 39]
@@
In ViTO, the attention mechanism serves as a direct parameterization of the integral operator $\mathcal{K}$. The continuous integral operator is discretized and parameterized via the attention mechanism as follows:[87]
$\mathbf{z}_{i}=\sum_{j=1}^{N}A_{ij}\mathbf{v}_{j}\approx\int_{D}\kappa\Big(x,y,v_{l}(x),v_{l}(y);\theta\Big)v_{l}(y)dy,\quad\forall x\in D.$ (13)
@@
Based on this operator definition, ViTO architecture adopts an explicit iterative strategy. The model consists of a stack of $L$ Transformer layers, where each layer $l$ possesses a distinct set of learnable parameters $\theta_{l}$. The update rule for the feature field from layer $l$ to $l+1$ is given by:[87]
@@
Although stacking independent Transformer modules enhances the model’s depth and expressive power, it inevitably introduces parameter redundancy. To further enhance parameter efficiency, we introduced an implicit variant of the model, termed IViTO. The architecture of the IViTO is illustrated in Fig. 2. Unlike stacking $L$ standard blocks with independent parameters as shown in Fig. 1(a), IViTO employs a weight-sharing mechanism that iteratively applies the same vision Transformer operator.[15, 1] Its iteration process can be written as:[87]
@@
II.3 Physics-Informed Transformer operator
@@
$\begin{split}\mathcal{L}_{1}&=\|\nabla\cdot\bar{\mathbf{u}}\|_{L^{2}(T;D)}^{2},\\ \mathcal{L}_{2}&=\|\partial_{t}\bar{\mathbf{u}}+\bar{\mathbf{u}}\cdot\nabla\bar{\mathbf{u}}+\nabla\bar{p}-\nu\nabla^{2}\bar{\mathbf{u}}-\nabla\cdot\bm{\tau}\|_{L^{2}(T;D)}^{2},\\ \mathcal{L}_{PDE}&=\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{2},\end{split}$ (16)
@@
In numerical simulations of statistically steady-state turbulence, the external forcing term $\mathcal{F}$ is typically implemented through spectral space manipulations, scaling velocity modes in the low-wavenumber region to preserve the target energy spectrum.[17] Let $\hat{\mathbf{u}}$ denote the velocity field in spectral space. We define rescaling operator $\mathcal{R}_{f}$ formulated as:[96]
$\hat{\mathbf{u}}^{f}(k)=\alpha(k)\hat{\mathbf{u}}(k),\quad\text{where }\alpha(k)=\begin{cases}\sqrt{E_{0}(1)/E(1)},&0.5\leq k\leq 1.5,\\ \sqrt{E_{0}(2)/E(2)},&1.5\leq k\leq 2.5,\\ 1,&\text{otherwise}.\end{cases}$ (18)
@@
While physics-informed Transformer models have rarely been applied to three-dimensional problems, the physics-informed Fourier neural operator (PIFNO) has demonstrated a effective performance for 3D turbulence in our previous study.[96] The implementation details of the surrogate model FNO are provided in Appendix A.
III Numerical results

### v4

@@ Abstract
keywords:
Physics-Informed Neural Operator, Transformer, Turbulence, Large-Eddy Simulation
1 Introduction
[…]
@@ 1 Introduction
In recent years, deep neural network methods have been developed for directly simulating flow fields [17, 18, 19, 20, 21]. Raissi et al. proposed physics-informed neural networks (PINNs), a method that embeds physical laws directly into the loss function to address forward and inverse partial differential equations (PDE) problems [22]. Wang et al. proposed the turbulent flow net (TF-Net), a physics-informed architecture designed to capture the dynamics of two-dimensional (2D) Rayleigh-Bénard convection [23]. Jin et al. developed NSFnets, a PINN-based framework that predicts flow fields given their initial and boundary conditions [24]. Although PINNs have demonstrated the capability to achieve satisfactory results even with sparse or no labeled data by embedding physical laws into the loss function, their application in complex scenarios is a great challenge [25, 26]. Wang et al. proposed an adaptive weighting algorithm to mitigate gradient pathologies in PINNs by dynamically balancing the gradient magnitudes of different loss terms [27]. Jagtap et al. proposed adaptive activation functions to accelerate the convergence and improve the accuracy of PINNs by dynamically adjusting the slopes of neurons during training [28]. Yang et al. introduced Bayesian physics-informed neural networks (B-PINNs), which combine Bayesian inference with PINNs to enable stable predictions and uncertainty quantification of physical fields using noisy experimental data [29]. Wang et al. achieved the simulations of 3D turbulent flows by improving PINN training strategies [30]. Wang et al. introduced ShampoO with Adam in the Preconditioner’s eigenbasis (SOAP), a quasi-Newton method that resolves gradient conflicts through second-order optimization, achieving breakthrough accuracy in turbulence simulations at high Reynolds numbers [31]. Zhang et al. used coordinate transformation to mitigate large gradients, enabling accurate PINN simulations of high-Reynolds-number flat plate and NACA0012 airfoil flows [32], while Cao et al. utilized mesh transformation to resolve the sharp leading-edge transitions in subsonic airfoil flows [33]. Moreover, Cao et al. stabilized PINN training by introducing a Jacobian-based gain-tuning strategy, enabling the robust simulation of complex 3D flows [34]. To improve the temporal modeling of PINNs, Zhao et al. proposed Pinnsformer, which leverages Transformer blocks to better capture temporal dependencies than conventional MLPs [35].
@@
Previous studies primarily focused on neural networks for learning finite-dimensional mappings, often struggling to generalize to other flow types or boundary conditions. To address this limitation, the framework of neural operators (NO) was proposed to learn operators between infinite-dimensional function spaces [36, 37, 38, 39, 40, 41]. Among these, Li et al. proposed the Fourier neural operator (FNO), which approximates the solution operator by parameterizing the integral kernel in the Fourier domain. In 2D Navier-Stokes turbulence, FNO outperformed convolutional networks in terms of accuracy and efficiency [42]. Consequently, numerous extensions of FNO were proposed. You et al. proposed the implicit Fourier neural operator (IFNO), which significantly reduces memory requirements and the number of trainable parameters [43]. Tran et al. proposed the factorized Fourier neural operator (F-FNO), which achieves more efficient operator learning by performing low-rank factorization on the frequency-domain weight tensor [44]. Additionally, Peng et al. embedded an attention mechanism into the FNO model, enhancing its predictive accuracy for 2D turbulence [45]. While early FNO-based studies primarily focused on one-dimensional (1D) and 2D problems, Li et al. extended the framework to 3D space and applied it to the predictions of 3D homogeneous isotropic turbulence [46]. To further address the long-term instability issue in FNO, Li et al. subsequently developed an implicit U-Net enhanced FNO (IUFNO), achieving stable and accurate long-term predictions of 3D turbulent flows [47]. Wang et al. extended IUFNO to 3D turbulent channel flows, achieving fast, accurate and stable long-term predictions [48].
@@
Although learning parameters in the frequency domain enables FNO to predict turbulence effectively and accurately, it should be noted that FNO relies on the Fourier transform and is therefore primarily applicable to periodic boundaries [49]. Recently, Transformer models have demonstrated outstanding performance in computer vision and natural language processing [50, 51, 52, 53]. Similar to FNO, Transformer also has the ability to capture global information and is increasingly being employed to solve partial differential equations [54, 55, 56] and perform super-resolution flow reconstruction [57, 58, 59, 60, 61]. Cao et al. established a theoretical connection between self-attention and Petrov-Galerkin projection, and proposed Galerkin Transformer to enhance approximation efficiency in PDE problems [62]. Patil et al. proposed a convolutional Transformer model, achieving accurate long-term predictions for 2D wake flows [63]. Xu et al. proposed a super-resolution Transformer (SRTT) to recover fine-scale structures of turbulent flows from low-resolution data [64].
@@
Despite these successes, the high computational cost of standard attention remains a challenge for 3D simulations. To address this issue, Niu et al. proposed a 3D spatiotemporal model incorporating a channel attention mechanism to predict unsteady airflow over complex three-dimensional aircraft structures [65]. Li et al. proposed the factorized Transformer (FactFormer) for 3D PDE simulations via axial factorization [66]. Yang et al. achieved stable long-term predictions of 3D turbulence by proposing the implicit factorized Transformer (IFactFormer) [67]. Yang et al. subsequently improved the model by replacing serial attention with a parallel mechanism, which achieves more accurate long-term predictions of 3D turbulent channel flows [68].
@@
Another effective approach is the vision Transformer (ViT), proposed by Dosovitskiy et al., which reduces computational costs while achieving excellent results in computer vision by partitioning images into patches [53]. Recently, the ViT has been widely applied to fluid mechanics and aerodynamics problems [69, 70, 71, 72, 73, 74, 75, 76]. Deng et al. proposed a modified ViT-based network for the rapid and accurate predictions of transonic flow fields over supercritical airfoils [77]. Ovadia et al. proposed the vision Transformer neural operator (ViTO), which exhibits an excellent predictive accuracy in 2D partial differential equations [78]. Wang et al. proposed the continuous vision Transformer (CViT) for 2D PDEs [79].
@@
However, purely data-driven neural operators often suffer from reliance on high-fidelity training data and a lack of physical interpretability. Recently, the physics-informed method has been introduced into neural operators to solve PDEs, including physics-informed DeepONets [80] and physics-informed Fourier neural operator [81]. It has been shown that physics-informed neural operator (PINO) models perform with high accuracy across various differential equations [82, 83, 84, 85, 86, 87, 88, 89]. By leveraging equation constraints, PINO can learn partial differential equations with minimal or even no data. Zhao et al. proposed the large-eddy simulation nets model (LESnets), which performs excellently in predictions of 3D free shear turbulence and isotropic decaying turbulence without using any labeled data [90]. Recently, Lorsung et al. proposed the physics informed token Transformer (PITT) to solve 1D and 2D PDEs [91]. Due to the increase in sequence length, applying Transformer methods to 3D problems requires significant computational costs. Consequently, we consider the ViT method as an effective solution to handle the computational complexity of 3D problems.
@@
2 Methodology
2.1 Governing equations of large eddy simulation
The Navier-Stokes equations governing incompressible fluid motion are given by [1]:
@@ 2.1 Governing equations of large eddy simulation
Due to the multi-scale nature of turbulence at high Reynolds number, directly resolving all spatial scales requires extremely fine grids, resulting in computational costs that are expensive for practical engineering applications. The LES method applies a spatial filter to the Navier-Stokes equations, enabling the use of coarser grids to resolve large-scale turbulent structures while modeling the unresolved small-scale motions. The filtered governing equations can be written as [1]:
@@
A widely used approach to close the SGS stress is the Smagorinsky (SM) model, which is a representative functional model based on the eddy-viscosity hypothesis [4, 92]. The SM model assumes that the SGS stress is proportional to the resolved strain rate tensor $\bar{S}_{ij}$, formulated as:
@@
where $C_{\text{smag}}$ is the Smagorinsky coefficient and $\Delta$ is the filter width. $\bar{S}_{ij}=\frac{1}{2}\left(\frac{\partial\bar{u}_{i}}{\partial x_{j}}+\frac{\partial\bar{u}_{j}}{\partial x_{i}}\right)$ denotes the filtered rate-of-strain tensor. $|\bar{S}|=\sqrt{2\bar{S}_{ij}\bar{S}_{ij}}$ represents the characteristic strain rate.
The integral length scale and the large-eddy turnover time are defined as [1]:
@@
2.2 Vision Transformer operator
In recent years, Transformer [50] based on self-attention has become the preferred models in natural language processing (NLP). However, it is computationally prohibitive to apply self-attention directly to images due to the quadratic cost associated with the number of pixels. To address this issue, Dosovitskiy et al. proposed the vision Transformer (ViT) [53]. ViT reduces the sequence length by partitioning images into patches, thereby reducing the computational overhead of self-attention. After extensive data pre-training, it achieved superior results compared to convolutional neural networks (CNNs).
@@ 2.2 Vision Transformer operator
To approximate the nonlinear solutions operator $\mathcal{G}$, the network is constructed as an iterative architecture. Each layer updates the feature field $v_{l}$ via a kernel integral operator $\mathcal{K}$, defined generally as [90, 47, 42]:
@@
In ViTO, the attention mechanism serves as a direct parameterization of the integral operator $\mathcal{K}$. The continuous integral operator is discretized and parameterized via the attention mechanism as follows [68]:
$\begin{split}\mathbf{z}_{i}&=\sum_{j=1}^{N}A_{ij}\mathbf{v}_{j}\\ &\approx\int_{D}\kappa\Big(x,y,v_{l}(x),v_{l}(y);\theta\Big)v_{l}(y)dy,\quad\forall x\in D.\end{split}$ (13)
@@
Based on this operator definition, ViTO architecture adopts an explicit iterative strategy. The model consists of a stack of $L$ Transformer layers, where each layer $l$ possesses a distinct set of learnable parameters $\theta_{l}$. The update rule for the feature field from layer $l$ to $l+1$ is given by [68]:
@@
Although stacking independent Transformer modules enhances the model’s depth and expressive power, it inevitably introduces parameter redundancy. To further enhance parameter efficiency, we introduced an implicit variant of the model, termed IViTO. The architecture of the IViTO is illustrated in Fig. 2. Unlike stacking $L$ standard blocks with independent parameters as shown in Fig. 1(a), IViTO employs a weight-sharing mechanism that iteratively applies the same vision Transformer operator [93, 94]. Its iteration process can be written as [68]:
@@
2.3 Physics-Informed Transformer operator
@@
$\begin{split}&\mathcal{L}_{1}=\|\nabla\cdot\bar{\mathbf{u}}\|_{L^{2}(T;D)}^{2},\\ &\mathcal{L}_{2}=\|\partial_{t}\bar{\mathbf{u}}+\bar{\mathbf{u}}\cdot\nabla\bar{\mathbf{u}}+\nabla\bar{p}-\nu\nabla^{2}\bar{\mathbf{u}}-\nabla\cdot\bm{\tau}\|_{L^{2}(T;D)}^{2},\\ &\mathcal{L}_{PDE}=\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{2},\end{split}$ (16)
@@
In numerical simulations of statistically steady-state turbulence, the external forcing term $\mathcal{F}$ is typically implemented through spectral space manipulations, scaling velocity modes in the low-wavenumber region to preserve the target energy spectrum [95]. Let $\hat{\mathbf{u}}$ denote the velocity field in spectral space. We define rescaling operator $\mathcal{R}_{f}$ formulated as [90]:
$\begin{split}&\hat{\mathbf{u}}^{f}(k)=\alpha(k)\hat{\mathbf{u}}(k),\\ &\text{where }\alpha(k)=\begin{cases}\sqrt{E_{0}(1)/E(1)},&0.5\leq k\leq 1.5,\\ \sqrt{E_{0}(2)/E(2)},&1.5\leq k\leq 2.5,\\ 1,&\text{otherwise}.\end{cases}\end{split}$ (18)
@@
While physics-informed Transformer models have rarely been applied to three-dimensional problems, the physics-informed Fourier neural operator (PIFNO) has demonstrated a effective performance for 3D turbulence in our previous study [90]. The implementation details of the surrogate model FNO are provided in Appendix A.
3 Numerical results